OpenStudy (anonymous):

Use the distributive property to simplify the expression. -6(1^2+3)-2(1^2-2)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i am going to guess that there is a typo there very rarely do you see \(1^2\)

OpenStudy (misty1212):

\[-6(1^2+3)-2(1^2-2)\]is a number no need for the distributive law could it be \[-6(x^2+3)-2(x^2-2)\]or something like that?

OpenStudy (anonymous):

yeah it was like that but it wouldn't let me put it like that

OpenStudy (misty1212):

with the \(x\)?

OpenStudy (anonymous):

it doesn't say

OpenStudy (anonymous):

it just say t^2

OpenStudy (misty1212):

can you post a screen shot or an attachment?

OpenStudy (misty1212):

ooh it is a \(t\) ok

OpenStudy (misty1212):

\[-6(t^2+3)-2(t^2-2)\] right?

OpenStudy (anonymous):

-6(t^2+3)-2(t^2-2)

OpenStudy (misty1212):

ok multiply each term in the first parentheses by \(-6\) what do you get?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

I think I calculated it wrong

OpenStudy (misty1212):

\[-6(t^2+3)=-6\times t^2+-6\times 3\] is a start